### v1

# Active Alignments of Lens Systems with Reinforcement Learning*

Abstract
Aligning a lens system relative to an imager is a critical challenge in camera manufacturing. While optimal alignment can be mathematically computed under ideal conditions, real-world deviations caused by manufacturing tolerances often render this approach impractical. Measuring these tolerances can be costly or even infeasible, and neglecting them may result in suboptimal alignments. We propose a reinforcement learning (RL) approach that learns exclusively in the pixel space of the sensor output, eliminating the need to develop expert-designed alignment concepts. We conduct an extensive benchmark study and show that our approach surpasses other methods in speed, precision, and robustness. We further introduce relign, a realistic, freely explorable, open-source simulation utilizing physically based rendering that models optical systems with non-deterministic manufacturing tolerances and noise in robotic alignment movement. It provides an interface to popular machine learning frameworks, enabling seamless experimentation and development. Our work highlights the potential of RL in a manufacturing environment to enhance efficiency of optical alignments while minimizing the need for manual intervention.
I Introduction
The assembly of optical devices requires precise positioning when joining their individual components. This requirement is essential in a wide range of products, including cameras in mobile phones, fiber optics, aerial and medical imaging and optical projection systems for microlithography [1]. One particularly sensitive process is an alignment, where two components must be precisely positioned relative to each other to achieve high precision. A prominent example is the positioning of a lens system relative to an imager [2]. The compound product must be assembled in a way that the optical performance is maximized. While high-cost lenses are often designed to ease the alignment with an imager, achieving optimal alignment with low-cost components presents a significant challenge. All of these components typically offer many degrees of freedom, each influencing multiple performance metrics in complex and interdependent ways. Often, it is unclear how the position has to be modified in order to reach a performance satisfying predefined quality constraints. Additionally, variations of the components make the relations between the position and the optical performance diffuse and noisy. These challenges have been widely studied in the literature [3, 4, 5, 6].
[…]
Figure 1: Schematic presentation of a single alignment step, where a lens system consisting of three single lenses (left) has to be positioned relative to an optical sensor (right).
The classic way to deal with such problems involves extensive scans during the alignment of each optical system individually, where any degrees of freedom are varied and evaluated separately. To make those scans robust against manufacturing tolerances within the components, many possible positions are scanned by a structured walk through the alignment space, for example along coordinate axes. Often, these algorithms solely rely on hand-crafted features obtained from the high-dimensional sensor output, where sensor and movement noise make it hard to conduct deterministic algorithms. Speeding up optical alignments has thus been a fruitful application of machine learning methods in the past, see [7] for a review. Some approaches predict next alignment moves from misaligned settings in a supervised fashion [8, 9]. A detailed study for Fast-Axis Collimating Lenses can be found in [10]. In their basic form, however, active alignment problems are no supervised learning problems. This is due to the fact that first, symmetries and offsets in the optical layout make it hard—or even impossible—to set up a supervised dataset from the sensor observation to the optimal sensor image. Second, training models via supervised learning cannot account to minimize the number of alignment steps. For instance, sometimes a step into the wrong direction has to be taken in order to explore symmetries in the robotic movements. Thus, more naturally, optical alignments are modeled as an RL problem which canonically allows training models to find short trajectories to optimal positions. RL algorithms have demonstrated the ability to learn complex relationships for various challenging tasks [11, 12, 13]. There has also been plenty of research using RL for process control in manufacturing (see [14] and references therein). For the alignment of laser optics or interferometers, RL has already been applied successfully as demonstrated in [15] or [16], respectively. Particularly when applied to real systems, RL comes with its own intrinsic challenges [17], like sparse and delayed rewards [18], data inefficiencies [19], and reproducibility issues [20], rendering high need for research when applied to new tasks.
In this study, we formulate optical alignment problems as an RL task, where optimal robotic alignment movements are learned solely in the pixel space from a high-dimensional sensor observation (see Figure 1). The alignment goal is reached when the difference between the observed image and a given reference pattern falls below a predefined threshold. We study different reward functions to motivate RL agents to find optimal alignment positions in as few steps as possible. To the best of our knowledge, our work is the first that treats an optical alignment task as a Partially Observable Markov Decision Process (POMDP). More specifically, our main contributions are:
  • We formulate active lens alignment problems as POMDP tangible by RL algorithms.
  • We introduce the Python framework relign [^1] Code under https://github.com/hs-kempten/relign. for simulating active alignment scenarios using the physically based rendering framework Mitsuba [21].
  • We provide an interface compatible with the Gymnasium API [22] for benchmarking state-of-the-art methods on representative alignment tasks, including RL models and black-box optimization techniques.
  • We show that RL algorithms can solve real-world-inspired alignment problems more efficiently than other methods based on machine learning, even under presence of manufacturing tolerances and noise in robotic alignment movements while maintaining high accuracy and low inference time.
Our approach not only speeds up the optical alignment process significantly, it also renders the need to design hand-crafted features obsolete.
II Active Alignments of Lens Systems
II-A Problem Formulation
An optical alignment process joins multiple components to maximize optical performance. In this work, the component to be aligned is a lens system consisting of a fixed number of single lenses $L=(L_{1},\ldots,L_{k})$ that need to be positioned relative to an optical sensor. This situation is typical when manufacturing cameras. The goal of an alignment is to move $L$ to a position $s=(x,y,z,R_{x},R_{y},R_{z})\in\mathbb{R}^{6}$ relative to a sensor, typically with an automated robotic alignment system, such that the optical performance is maximal.
The main challenge is that in each alignment, the process has to adjust to new conditions, mainly due to variances within $L$ and when gripping $L$. The first type of variances is a randomized offset arising from the placement of $L$ in the alignment station. This can be modeled by a randomized starting position $W_{\mathrm{off}}\in\mathbb{R}^{6}$ representing a random translation and rotation offset. When connecting to the robotic alignment system, the movement is typically not optimal, meaning that when a movement $a\in\mathbb{R}^{6}$ in a state $s$ is executed, the new state is not $s+a$ but the slightly distorted state $s+W_{\mathrm{dist}}\cdot a$ with $W_{\mathrm{dist}}\in\mathbb{R}^{6\times 6}$. Moreover, the new state can be clipped, for instance because a boundary condition is met. We refer to Section Section IV-B for more details on how $W_{\mathrm{dist}}$ is constructured and how the boundary behavior is modeled. Another type of variances comes from the production of $L$ itself, which cannot be changed by the process during the alignment. That is, each single lens $L_{i}$ in $L$ has an individual tilt and position offset in comparison to the ideal lens system (see Figure 2). We denote the offsets of each lens within $L$ by $W_{\mathrm{L}}\in\mathbb{R}^{k\times 6}$. Other variances not considered in this work are dispersions in the geometries of the single lenses, like their curvature. Here, the variances $W=(W_{\mathrm{off}},W_{\mathrm{dist}},W_{\mathrm{L}})$ characterize an alignment completely and we assume that these are sampled from an unknown distribution $W\sim\rho$. We further assume that the variances $W$ are latent to the alignment process, i.e., cannot be measured while aligning. As a result, it is not possible to directly compute the position where optical performance is optimal using the physical equations of e.g. geometrical optics.
Refer to caption Refer to caption $W_{\mathrm{L}}=0$$W_{\mathrm{L}}\neq 0$
Figure 2: Impact of $W_{\mathrm{L}}$ on lenses. Here, the arrows depict normal vectors of the lenses to visualize noise in the rotations.
II-B Performance Measurements
To quantify the optical performance at a given state $s$, collimated light is sent through $L$ and measured at the sensor with width $w$ and height $h$, yielding a high-dimensional image $O_{W}(s)\in\mathbb{R}^{w\times h}$ of light intensities. The sensor output $O_{W}(s)$ is noisy, and retrieving $O_{W}(s)$ multiple times for the same position $s$ always yields slightly different observations. We call this noise sensor variance. In many industrial applications, hand-crafted scalar features are extracted from $O_{W}(s)$, typically involving the optical transfer functions, for which quality bounds have to be reached during the alignment. Here, however, we study problems where a generic reference output $O^{*}$ independent of $W$ is given such that the alignment task is the identification of a state $s^{*}$ where the sensor output matches the reference, i.e., $O_{W}(s^{*})\approx O^{*}$. The reference pattern can be considered as $O^{*}=\mathbb{E}[O_{W^{*}}(s^{*})]$, that is the sensor output without noise $W^{*}=0$ in the positioning of $L$ and its lenses at the optimal position $s^{*}$. In practice, the choice of the light field used for creating a test pattern at sensor level is dictated by the specific application requirements, such as whether the camera needs to achieve sharp focus in the near or far field. A common example is the Siemens star [23], which is widely used for evaluating optical performance of digital cameras [24]. Numerous methods exist to measure the quality of an optical image based on projected patterns such as a pattern of Siemens stars and slanted edges as described in ISO 12233 [24]. Given a reference pattern, the optimization problem
$\operatorname*{arg\,min}_{s\in S}\|O_{W}(s)-O^{*}\|$ (1)
[…]
Figure 3: Visualization of $(s_{i},s_{j})\to\min_{s_{k},k\neq i,k\neq j}\|O_{W}(s)-O^{*}\|$ for each tuple $\{i,j\}\subset\{x,y,z,R_{x},R_{y}\}$.
II-C Alignment Algorithms
As the variances $W$ are unobservable by the alignment process, most active alignment algorithms explore the state space $S$ iteratively by selecting states and searching for directions where the score $\|O_{W}(s)-O^{*}\|$ attains its minimum. Once positioned at a new state, a new sensor signal is observed. An alignment algorithm computes a series of subsequent actions $a_{1},\ldots,a_{n}$, step by step, starting from a randomized initial position $s_{0}=W_{\mathrm{off}}$. Each action is an element of a set of allowed actions $A\subset\mathbb{R}^{6}$ and generates a sequence of states $s_{i}=s_{0}+W_{\mathrm{dist}}(a_{1}+\ldots+a_{i})$ such that $s_{n}$ is optimal, i.e., $\|O_{w}(s_{n})-O^{*}\|\leq\theta$. The computation of $a_{i}$ must be based on a subset of the $i-1$ many images $O_{W}(s_{1}),\ldots,O_{W}(s_{i-1})$ obtained so far. In Section V-B, we state the alignment algorithms used in our benchmark study.
III Approach
III-A Alignments as a POMDP
In this section, we describe how RL algorithms can be used as alignment algorithms as defined in Section II-C. Specifically, we consider active alignments as an episodic POMDP, where each episode is the alignment of a given lens system $L$. As defined in Section II, $S$ represents the set of states, and $A$ denotes the set of actions. In the alignment of $L$, including variances $W=(W_{\mathrm{off}},W_{\mathrm{dist}},W_{\mathrm{L}})$, selecting $a$ at $s$ results in the new state $s^{\prime}=s+W_{\mathrm{dist}}\cdot a$, yielding a reward $R(s,s^{\prime})$. A detailed view on different reward functions used is provided in Section III-B. The state $s\in S$ cannot be directly observed. Instead, only the high-dimensional sensor output $O_{W}(s)$ is given, which is controlled by a conditional probability density function depending on $s$. We explain in Section IV-A how an image $O_{W}(s)$ is sampled from the sensor at a given state $s$. For given $W$, an episode ends once a terminal state, that is an optimal state from $\{s\in S:\|O_{W}(s)-O^{*}\|\leq\theta\}$, or an upper limit of steps $T\in\mathbb{N}$ is reached. The goal is to find a policy $\pi$ that maps observations to actions in a way that maximizes the accumulated observed reward. More formally, given the observation $O_{W}(s)$ for $s$, the next state is $s+W_{\mathrm{dist}}\cdot\pi(O_{W}(s))$. Starting from an initial state $s_{0}$, this combined dynamics of sampled observations from the sensor and generated actions by the policy $\pi$ yields a trajectory $(s_{0},\ldots,s_{k})$ of subsequent states with $s_{i+1}=s_{i}+W_{\mathrm{dist}}\cdot\pi(O_{W}(s_{i}))$ where the last state $s_{k}$ is either optimal or $k=T$ holds. The goal is to find a policy such that
$\mathbb{E}_{W\sim\rho,(s_{0},\ldots,s_{k})\sim\pi}\left[\sum_{i=0}^{k-1}\gamma^{i}R(s_{i},s_{i+1})\right]$ (2)
is maximized, where $\gamma\in(0,1)$ is the discount factor given to trade-off rewards in early and late states. Note that in general the reward also depends on the action taken, which is not required in this work and hence omitted.
III-B Rewards
To train effective agents for an alignment task, the reward function $R$ has to be designed in a way that the optimization task from Equation (2) is solved in a minimal number of steps. We consider two distinct reward functions: The first, $R_{\mathrm{step}}$, assigns a fixed penalty of $-1$ for each step taken, independent of the states and their distance to the goal, i.e., $R_{\mathrm{step}}(s,s^{\prime})=-1$. The second approach involves rewarding each step based on the extent to which the action brings the system closer to the optimal state. More formally, we define
$R_{\mathrm{pot}}(s,s^{\prime})=\|O_{W}(s)-O^{*}\|-\|O_{W}(s^{\prime})-O^{*}\|.$
That is, a larger reward is obtained if $s^{\prime}$ is closer to the optimal state than $s$. Note that $R_{\mathrm{pot}}$ is a potential-based reward [25], that means that if $\gamma=1$, the reward of a trajectory $(s_{0},\ldots,s_{k})$ satisfies $\sum_{i=0}^{k-1}R_{\mathrm{pot}}(s_{i},s_{i+1})=R_{\mathrm{pot}}(s_{0},s_{k})$.
IV Simulating Alignments
One main challenge in simulating realistic optical alignments is to accurately model how a sensor measures light emitted from a source and propagated through optical lenses. Here, we use Mitsuba3, a physically based rendering engine for forward and inverse light-transport simulation. This not only allows calculating light intensities $O_{W}(s)$ measured at the sensor (see Section IV-A) for a concrete position $s$, but also changing $s$ dynamically (see Section IV-B). All Mitsuba scenes consist of a sensor, $k$ many identical biconvex single lenses, and a light source.
IV-A Sensor Outputs
The Mitsuba scene emulates collimated light using a binary environment map with a centered rectangular light source. Starting at the sensor, Mitsuba traces light rays backwards that pass through the lenses. We use an irradiance meter as sensor, which measures the incident power per unit area over a predefined shape. In our setup, the sensor shape is $0.5\times 0.5$ in Mitsuba space coordinates. To solve the high-dimensional problem of rendering $O_{W}(s)$ numerically, Mitsuba employs Monte Carlo integration, drawing $s_{c}$ many samples from a uniform distribution. As a consequence, the sensor output $O_{W}(s)$ is different when rendering the same position $s$ with different seeds. This leads to the probability density function of the POMDP as described in Section III-A. An example of the resulting sensor output is shown in Figure 1. The measurements are interpreted as a grayscale image and stored without any post-processing.
IV-B Position Changes
Without loss of generality, we assume that the set of all possible states is the unit interval $[0,1]^{6}$. We place $k$ many lenses in the lens system $L$ in a way that $s^{*}:=(0.5,\ldots,0.5)\in\mathbb{R}^{6}$ is the optimal position. Upon initialization, we sample $W_{\mathrm{L}}$ from a normal distribution and reposition the single lenses within $L$ accordingly. Afterward, a starting vector $W_{\mathrm{off}}\in[0,1]^{6}$ is sampled uniformly which defines the initial state $s_{0}:=W_{\mathrm{off}}$. The movement distortion matrix is constructed as $W_{\mathrm{dist}}=I_{6}+{\epsilon}\in\mathbb{R}^{6\times 6}$ where each coordinate in ${\epsilon}\in\mathbb{R}^{6\times 6}$ is sampled from a normal distribution. The positioning of the lens system can be varied by setting an action $a\in\mathbb{R}^{6}$ leading to an update $s^{\prime}=s+W_{\mathrm{dist}}\cdot a$. To ensure that $s^{\prime}$ stays within $[0,1]^{6}$, each coordinate of $s+W_{\mathrm{dist}}\cdot a$ is clipped into $[0,1]$ before updating the Mitsuba scene. Subsequently, the sensor output $O_{W}(s^{\prime})$ is generated as described in Section IV-A.
IV-C Generating the Reference Pattern
To decide whether a state $s$ is in the optimality condition, the observation $O_{w}(s)$ has to be compared to a reference pattern $O^{*}$. This reference pattern only depends on the number of lenses $k$ in the lens system, not on the noise level of $W$. To generate $W$ in our synthetic setting, we use a perfectly aligned lens system where each of the $k$ many lenses are perfectly aligned as well, i.e., $W^{*}:=0$. We then sample $1.000$ observations from $O_{W^{*}}(s^{*})$ and compute the pixel-wise mean image as an approximation for $O^{*}=\mathbb{E}[O_{W^{*}}(s^{*})]$.
V Experiments
V-A Benchmark Environments
When rendering images, one typically faces a trade-off between computational efficiency and image quality. To reduce overall RL training time, we prioritize fast image generation at the cost of lower pixel resolution and increased sensor variance. Because of that, we only use $50\times 50$ pixels and $512$ samples per pixel for rendering. This also motivates the decision to use a simple rectangle as reference pattern, as more detailed patterns are not recognizable at this resolution.
For the evaluation, we focus on six distinct benchmark setups. Each setup considers a set of $k\in\{2,3\}$ single lenses with radii of curvature of $-1.0$ and $3.8$ and thickness of $0.2$. Each single lens can have none, low or high individual variances $W_{L}$. The first three components of $W_{L}$ represent translation offsets along the $x$, $y$ and $z$ axes and are sampled from a normal distribution $(W_{L})_{i}\sim N_{0,6.25}$ for low and $(W_{L})_{i}\sim N_{0,12.5}$ for high variances with $i\in\{1,2,3\}$. The remaining two represent rotation offsets along the $x$ and $y$ axes, sampled as $(W_{L})_{j}\sim N_{0,0.00625}$ for low and $(W_{L})_{j}\sim N_{0,0.0125}$ for high variances with $j\in\{4,5\}$. These variances account for potential manufacturing imperfections within a certain tolerance bound. To create a challenging alignment situation, they are intentionally set very high (see Figure 2 for an extreme case). Furthermore, rotation around $z$ is considered redundant, as in a perfectly aligned scenario, rotation around the $z$ axis has no effect. Note that $O^{*}$ does not change for setups including object variances.
V-B Algorithms
Expert-designed alignment algorithms typically rely on hand-crafted features extracted from $O_{W}(s)$ that need to be optimized simultaneously. These features as well as their optimality region highly depend on the exact use case. To have a fair and generic comparison of our RL approach with conventional learning methods, we decided to benchmark against state-of-the art black-box optimizers to solve Equation (1) actively. For our study, we focused on three different models to solve the task of optical alignments:
V-B1 Proximal Policy Optimization
Proximal policy optimization (PPO) is a policy-based RL algorithm that trains a stochastic policy in an on-policy way [26]. To process the images, a CNN is used as policy network. While PPO is a versatile learning algorithm, many learning tasks are too complex to solve from scratch via RL. Instead of training an agent on a hard task from the start, curriculum learning structures the training by introducing simpler subtasks first, enabling the agent to build foundational skills before tackling more challenging scenarios. This typically leads to faster convergence, better generalization, and improved sample efficiency, especially in environments with sparse rewards or long-horizon decision-making problems [27]. The same holds true when training an RL agent for a too low optimality threshold $\theta$: The agent fails to reach the optimality region, even when allowing long trajectory lengths before truncation. Thus, we apply a curriculum strategy where $\theta$ is decreased gradually in an exponential fashion, that is, starting from an initial threshold $\theta_{0}$, the threshold of the $i$-th task is $\theta_{i}=C\cdot\theta_{i-1}$ with $C<1$ (in our experiments, we set $C=0.9$). We denote the task as solved when the agent manages to reach the optimality region defined by $\theta_{i}$ within $20$ steps in five subsequent online evaluations. The maximal length of an episode $T$ is set to $100$. All models have been trained for $1.5\cdot 10^{6}$ global steps with an initial learning rate of $1e^{-4}$ which is decreased by a factor of $0.8$ if the reward stagnates and which is reset to the initial value if the task of the curriculum is solved. Figure 4 shows the evolution of the reward along the curriculum for $k=3$ lenses. All trainings are executed on NVIDIA H100 GPUs requiring roughly $0.32$ seconds per global step resulting in a total train time of approximately five days.
V-B2 Bayesian Optimization
Bayesian Optimization (BO) is a strategy for minimizing functions $f$ that are expensive to evaluate by building a probabilistic model of the objective function [28, 29]. It selects new evaluation points by balancing exploration and exploitation using an acquisition function, which predicts where the function is likely to improve. This approach is particularly useful when function evaluations are costly, as it finds optimal solutions with relatively few evaluations. Setting $f_{W}(s):=\|O_{W}(s)-O^{*}\|$, an alignment problem can be interpreted as a black-box optimization problem. We evaluated BO algorithms using a different probabilistic model than implemented in scikit-optimize [30]: Gaussian Processes (BO-GP) [31] and Random Forests (BO-RF) [32]. As vanilla BO algorithms explore $f_{W}$ for each $W$ and without a priori information, the state space has to be explored first randomly costing unnecessary steps. Thus, we also tested the method proposed in [33] (TransferGP) that allows pre-training a Gaussian process on samples from problem instances $f_{W_{1}},\ldots,f_{W_{m}}$.
V-B3 Random
As a baseline, we implemented an algorithm that samples uniformly at random from the alignment space.
V-C Results
For evaluation, each algorithm was executed on $100$ different environments for each of the six benchmark situations as described in Section V-A. We compared approaches that operate without any a priori information and require no domain knowledge. Except for the RL algorithms, the algorithms used do not involve a training phase. To address this imbalance, we reduced the search space for baseline algorithms to approximately eight percent of the search space used for RL. Figure 5 shows the RL-based method trained with $R_{\mathrm{pot}}$ surpases in all scenarios all other algorithms in terms of convergence speed. Moreover, independent of the number of lenses and the noise level, RL-PPO reaches the optimality region in under ten steps. As expected, the benchmarks with larger noise levels converge at a higher error, because due to the lens variances, even the best alignment cannot reach the minimal render variance error. In a direct comparison, the potential-based reward $R_{\mathrm{pot}}$ converges slighly faster than $R_{\mathrm{step}}$ to the optimal region (see Figure 6).
Figure 4: Achieved reward $R_{\mathrm{pot}}$ during an online evaluation for $k=3$ lenses and different noise levels while optimality threshold $\theta$ is adjusted according to the curriculum.
Only considering the computation time, RL-PPO requires a constant 25 ms and BO-RF 95 ms per alignment step. In contrast, BO-GP becomes increasingly time consuming as the number of steps increases. For 20 steps, each step takes an average of 100 ms, while for 50 steps the time per step increases to 178 ms. For the pre-trained Gaussian processes of TransferGP, the processing time for each step increases with the number of instances it has been pre-trained on. Already when trained on $m=10$ instances with $100$ samples each, their processing time per step takes several minutes while their performance equals almost the performance of vanilla BO-GP. Due to their impractical computation time, we have not included pretrained GP models in our benchmark.
Figure 5: Comparison of alignment algorithms.
Figure 6: Evaluation of RMSE for the best PPO models. The upper plot compares two reward functions for a fixed number of single lenses (three). The lower plot compares two different lens systems using the best reward function.
VI Conclusion
This work introduced an RL approach for active alignments of optical components. Unlike traditional alignment methods that rely on expert-designed alignment concepts involving the computation of hand-crafted features, our approach learns optimal alignment strategies directly from high-dimensional sensor observations. By leveraging RL, we demonstrated that alignment tasks can be solved more efficiently, even in the presence of noise and manufacturing tolerances. However, the low inference time of RL-algorithms at runtime comes at the price of many training iterations. Our experiments show that RL-based alignment not only outperforms conventional machine learning approaches in terms of efficiency but also eliminates the need for manually designed features. This work opens the door for further exploration of RL in high-precision optical assembly, with potential applications in automated manufacturing, adaptive optics, and real-time calibration of complex optical systems. Future research could focus on improving sample efficiency, integrating domain adaptation techniques, and extending RL-based alignment to real-world hardware implementations.
ACKNOWLEDGMENT
We thank Marco F. Huber and our colleagues from the Institute of Machine Vision for their careful proofreading and constructive feedback on the manuscript.

### v2

@@
 […]
-Figure 1: Schematic presentation of a single alignment step, where a lens system consisting of three single lenses (left) has to be positioned relative to an optical sensor (right).
+Figure 1: Schematic presentation of a single alignment step, where a lens system consisting of four single lenses (left) has to be positioned relative to an optical sensor (right).
 The classic way to deal with such problems involves extensive scans during the alignment of each optical system individually, where any degrees of freedom are varied and evaluated separately.
@@
 More specifically, our main contributions are:
-  • We formulate active lens alignment problems as POMDP tangible by RL algorithms.
-  • We introduce the Python framework relign [^1] Code under https://github.com/hs-kempten/relign.
+  • We formulate general active alignment problems as POMDP tangible by RL algorithms. To the best of our knowledge, this is the first work in doing so.
+  • We introduce the modular Python framework relign [^1] Code under https://github.com/hs-kempten/relign.
 for simulating active alignment scenarios using the physically based rendering framework Mitsuba [21].
-  • We provide an interface compatible with the Gymnasium API [22] for benchmarking state-of-the-art methods on representative alignment tasks, including RL models and black-box optimization techniques.
+Its interface is compatible with the Gymnasium API [22] and allows effective benchmarking of state-of-the-art methods on representative alignment tasks.
   • We show that RL algorithms can solve real-world-inspired alignment problems more efficiently than other methods based on machine learning, even under presence of manufacturing tolerances and noise in robotic alignment movements while maintaining high accuracy and low inference time.
-Our approach not only speeds up the optical alignment process significantly, it also renders the need to design hand-crafted features obsolete.
+Finding accurate and comparable benchmarks for active alignments is hard, not only due to privacy constraints of companies maintaining them as part of their core know-how, but also due to the fact that optical products are very diverse and have different requirements.
+Here, our framework provides a first step towards a common benchmarking environment for active alignments.
+Moreover, we argue that using RL for alignments not only has the potential to significantly speed up the optical alignment process but also renders the need for designing hand-crafted features obsolete.
 II Active Alignments of Lens Systems
+Automated robotic alignment systems often rely on high-precision motion platforms such as hexapods or piezo stages, which provide movement in all six degrees of freedom.
+These mechanisms enable sub-micrometer adjustments, allowing the robot to finely correct tilt, shift, and focus errors until the desired alignment is reached.
+In this section, we formalize the active alignment problem and introduce the main challenges faced during the alignment process.
 II-A Problem Formulation
 An optical alignment process joins multiple components to maximize optical performance.
-In this work, the component to be aligned is a lens system consisting of a fixed number of single lenses $L=(L_{1},\ldots,L_{k})$ that need to be positioned relative to an optical sensor.
+In this work, the component to be aligned is a lens system consisting of a fixed number of single lenses $L=(L_{1},L_{2},L_{3},L_{4})$ that need to be positioned relative to an optical sensor.
 This situation is typical when manufacturing cameras.
@@
 The main challenge is that in each alignment, the process has to adjust to new conditions, mainly due to variances within $L$ and when gripping $L$.
-The first type of variances is a randomized offset arising from the placement of $L$ in the alignment station.
+The first type of variance is a randomized offset arising from the initial placement of $L$ in the alignment station.
 This can be modeled by a randomized starting position $W_{\mathrm{off}}\in\mathbb{R}^{6}$ representing a random translation and rotation offset.
@@
 Moreover, the new state can be clipped, for instance because a boundary condition is met.
-We refer to Section Section IV-B for more details on how $W_{\mathrm{dist}}$ is constructured and how the boundary behavior is modeled.
-Another type of variances comes from the production of $L$ itself, which cannot be changed by the process during the alignment.
-That is, each single lens $L_{i}$ in $L$ has an individual tilt and position offset in comparison to the ideal lens system (see Figure 2).
+We refer to Section IV-B for more details on how $W_{\mathrm{dist}}$ is constructed and how the boundary behavior is modeled.
+Another type of variance comes from the production of $L$ itself, which cannot be changed by the process during the alignment.
+That is, each single lens $L_{i}$ in $L$ has an individual tilt and position offset in comparison to the ideal lens system.
 We denote the offsets of each lens within $L$ by $W_{\mathrm{L}}\in\mathbb{R}^{k\times 6}$.
@@
 geometrical optics.
-Refer to caption Refer to caption $W_{\mathrm{L}}=0$$W_{\mathrm{L}}\neq 0$
-Figure 2: Impact of $W_{\mathrm{L}}$ on lenses. Here, the arrows depict normal vectors of the lenses to visualize noise in the rotations.
 II-B Performance Measurements
-To quantify the optical performance at a given state $s$, collimated light is sent through $L$ and measured at the sensor with width $w$ and height $h$, yielding a high-dimensional image $O_{W}(s)\in\mathbb{R}^{w\times h}$ of light intensities.
-The sensor output $O_{W}(s)$ is noisy, and retrieving $O_{W}(s)$ multiple times for the same position $s$ always yields slightly different observations.
+To quantify the optical performance at a given state $s$, collimated light is sent through $L$ and measured at the sensor with width $w$ and height $h$, yielding a high-dimensional image $O(s,W)\in\mathbb{R}^{w\times h}$ of light intensities.
+The sensor output $O(s,W)$ is noisy, and retrieving $O(s,W)$ multiple times for the same position $s$ always yields slightly different observations.
 We call this noise sensor variance.
-In many industrial applications, hand-crafted scalar features are extracted from $O_{W}(s)$, typically involving the optical transfer functions, for which quality bounds have to be reached during the alignment.
-Here, however, we study problems where a generic reference output $O^{*}$ independent of $W$ is given such that the alignment task is the identification of a state $s^{*}$ where the sensor output matches the reference, i.e., $O_{W}(s^{*})\approx O^{*}$.
-The reference pattern can be considered as $O^{*}=\mathbb{E}[O_{W^{*}}(s^{*})]$, that is the sensor output without noise $W^{*}=0$ in the positioning of $L$ and its lenses at the optimal position $s^{*}$.
+In many industrial applications, hand-crafted scalar features are extracted from $O(s,W)$, typically involving the optical transfer functions, for which quality bounds have to be reached during the alignment.
+Here, however, we study problems where a generic reference output $O^{*}$ independent of $W$ is given such that the alignment task is the identification of a state $s_{W}$ where the sensor output matches the reference, i.e., $O(s_{W},W)\approx O^{*}$.
+The reference pattern can be considered as $O^{*}=\mathbb{E}[O_{W}(s_{W})]$, that is the sensor output without noise $W^{*}=0$ in the positioning of $L$ and its lenses at the respective optimal position $s_{W}$.
 In practice, the choice of the light field used for creating a test pattern at sensor level is dictated by the specific application requirements, such as whether the camera needs to achieve sharp focus in the near or far field.
@@
 Given a reference pattern, the optimization problem
-$\operatorname*{arg\,min}_{s\in S}\|O_{W}(s)-O^{*}\|$ (1)
-[…]
-Figure 3: Visualization of $(s_{i},s_{j})\to\min_{s_{k},k\neq i,k\neq j}\|O_{W}(s)-O^{*}\|$ for each tuple $\{i,j\}\subset\{x,y,z,R_{x},R_{y}\}$.
+$\operatorname*{arg\,min}_{s\in S}\|O(s,W)-O^{*}\|$ (1)
+has to be solved for a given situation $W\sim\rho$, where $S\subset\mathbb{R}^{6}$ denotes the set of allowed positions.
+Here, $\|\cdot\|$ denotes a vector norm, like the Euclidean distance, and we interpret the input matrices as vectors.
+Figure 2 shows some 2D-projections of this optimization problem for the alignment situations introduced in Section VI.
+Typically, a threshold $\theta\in\mathbb{R}_{\geq 0}$ is given such that any $s$ that satisfies $\|O(s,W)-O^{*}\|\leq\theta$ is considered as optimal.
+However, as $s$ can only compensate positional offsets of $L$ given by $W_{\mathrm{off}}$ under distortions given by $W_{\mathrm{dist}}$ and not the variances $W_{\mathrm{L}}$ inside $L$, some lens systems may have $\|O(s,W)-O^{*}\|>\theta$ for all positions $s$.
+Figure 2: Visualization of the two-dimensional projection $(s_{i},s_{j})\mapsto\|O(s^{*}-s_{i}e_{i}-s_{j}e_{j},W^{*})-O^{*}\|$ for each tuple $\{i,j\}\subset\{x,y,z,R_{x},R_{y}\}$, that is, all except two dimensions are fixed to the optimal values.
 II-C Alignment Algorithms
-As the variances $W$ are unobservable by the alignment process, most active alignment algorithms explore the state space $S$ iteratively by selecting states and searching for directions where the score $\|O_{W}(s)-O^{*}\|$ attains its minimum.
+As the variances $W$ are unobservable by the alignment process, most active alignment algorithms explore the state space $S$ iteratively by selecting states and searching for directions where the score $\|O(s,W)-O^{*}\|$ attains its minimum.
 Once positioned at a new state, a new sensor signal is observed.
 An alignment algorithm computes a series of subsequent actions $a_{1},\ldots,a_{n}$, step by step, starting from a randomized initial position $s_{0}=W_{\mathrm{off}}$.
-Each action is an element of a set of allowed actions $A\subset\mathbb{R}^{6}$ and generates a sequence of states $s_{i}=s_{0}+W_{\mathrm{dist}}(a_{1}+\ldots+a_{i})$ such that $s_{n}$ is optimal, i.e., $\|O_{w}(s_{n})-O^{*}\|\leq\theta$.
-The computation of $a_{i}$ must be based on a subset of the $i-1$ many images $O_{W}(s_{1}),\ldots,O_{W}(s_{i-1})$ obtained so far.
-In Section V-B, we state the alignment algorithms used in our benchmark study.
-III Approach
-III-A Alignments as a POMDP
+Each action is an element of a set of allowed actions $A\subset\mathbb{R}^{6}$ and generates a sequence of states $s_{i}=s_{0}+W_{\mathrm{dist}}(a_{1}+\ldots+a_{i})$ such that $s_{n}$ is optimal, i.e., $\|O(s_{n},W)-O^{*}\|\leq\theta$.
+The computation of $a_{i}$ must be based on a subset of the i1i-1 many images $O(s_{1},W),\ldots,O(s_{i-1},W)$ obtained so far.
+In Section VI-A, we state the alignment algorithms used in our benchmark study.
+III Alignments as a POMDP
 In this section, we describe how RL algorithms can be used as alignment algorithms as defined in Section II-C.
 Specifically, we consider active alignments as an episodic POMDP, where each episode is the alignment of a given lens system $L$.
-As defined in Section II, $S$ represents the set of states, and $A$ denotes the set of actions.
-In the alignment of $L$, including variances $W=(W_{\mathrm{off}},W_{\mathrm{dist}},W_{\mathrm{L}})$, selecting $a$ at $s$ results in the new state $s^{\prime}=s+W_{\mathrm{dist}}\cdot a$, yielding a reward $R(s,s^{\prime})$.
-A detailed view on different reward functions used is provided in Section III-B.
-The state $s\in S$ cannot be directly observed.
-Instead, only the high-dimensional sensor output $O_{W}(s)$ is given, which is controlled by a conditional probability density function depending on $s$.
-We explain in Section IV-A how an image $O_{W}(s)$ is sampled from the sensor at a given state $s$.
-For given $W$, an episode ends once a terminal state, that is an optimal state from $\{s\in S:\|O_{W}(s)-O^{*}\|\leq\theta\}$, or an upper limit of steps $T\in\mathbb{N}$ is reached.
+Here, a state $(s,W)$ is decomposed in a position $s\in S$ and variances $W=(W_{\mathrm{off}},W_{\mathrm{dist}},W_{\mathrm{L}})$ as defined in Section II and the set $A$ denotes the set of actions.
+Selecting $a$ at $(s,W)$ results in the new state $(s^{\prime},W)$ with $s^{\prime}=s+W_{\mathrm{dist}}\cdot a$, yielding a reward $R(s,s^{\prime})$.
+As of the state representing the variances stays constant through an episode, this MDP is also an contextual MDP in the sense of [25, Section 3].
+The state $(s,W)$ cannot be directly observed.
+Instead, only the high-dimensional sensor output $O(s,W)$ is given, which is controlled by a conditional probability density function depending on $s$.
+We explain in Section IV-A how an image $O(s,W)$ is sampled from the sensor at a given state $s$.
+For given $W$, an episode ends once a terminal state, that is an optimal state from $\{s\in S:\|O(s,W)-O^{*}\|\leq\theta\}$, or an upper limit of steps $T\in\mathbb{N}$ is reached.
 The goal is to find a policy $\pi$ that maps observations to actions in a way that maximizes the accumulated observed reward.
-More formally, given the observation $O_{W}(s)$ for $s$, the next state is $s+W_{\mathrm{dist}}\cdot\pi(O_{W}(s))$.
-Starting from an initial state $s_{0}$, this combined dynamics of sampled observations from the sensor and generated actions by the policy $\pi$ yields a trajectory $(s_{0},\ldots,s_{k})$ of subsequent states with $s_{i+1}=s_{i}+W_{\mathrm{dist}}\cdot\pi(O_{W}(s_{i}))$ where the last state $s_{k}$ is either optimal or $k=T$ holds.
+More formally, given the observation $O(s,W)$ for $s$, the next state is $s+W_{\mathrm{dist}}\cdot\pi(O(s,W))$.
+Starting from an initial state $s_{0}$, this combined dynamics of sampled observations from the sensor and generated actions by the policy $\pi$ yields a trajectory $(s_{0},\ldots,s_{k})$ of subsequent states with $s_{i+1}=s_{i}+W_{\mathrm{dist}}\cdot\pi(O(s_{i},W))$ where the last state $s_{k}$ is either optimal or $k=T$ holds.
 The goal is to find a policy such that
@@
 Note that in general the reward also depends on the action taken, which is not required in this work and hence omitted.
-III-B Rewards
 To train effective agents for an alignment task, the reward function $R$ has to be designed in a way that the optimization task from Equation (2) is solved in a minimal number of steps.
-We consider two distinct reward functions: The first, $R_{\mathrm{step}}$, assigns a fixed penalty of $-1$ for each step taken, independent of the states and their distance to the goal, i.e., $R_{\mathrm{step}}(s,s^{\prime})=-1$.
-The second approach involves rewarding each step based on the extent to which the action brings the system closer to the optimal state.
-More formally, we define
-$R_{\mathrm{pot}}(s,s^{\prime})=\|O_{W}(s)-O^{*}\|-\|O_{W}(s^{\prime})-O^{*}\|.$
-That is, a larger reward is obtained if $s^{\prime}$ is closer to the optimal state than $s$.
-Note that $R_{\mathrm{pot}}$ is a potential-based reward [25], that means that if $\gamma=1$, the reward of a trajectory $(s_{0},\ldots,s_{k})$ satisfies $\sum_{i=0}^{k-1}R_{\mathrm{pot}}(s_{i},s_{i+1})=R_{\mathrm{pot}}(s_{0},s_{k})$.
+During training time of a optical system $L$ with variances $W$, agents can be provided with dense and insightful reward signals, for instance containing information of the distance to the desired optimum $s_{W}$, which need to be known beforehand.
+An ideal reward signal would be $\|s-s_{W}\|$ which, however, requires the knowledge of $s_{W}$.
+In our simulation and during train time, this could be computed, for instance using a conventional alignment algorithm, but this would slow down training time as this needs to be done for each training episode anew.
+Instead, we use an indirect signal which gives the distance to $s^{*}$, i.e.
+the optimal position if $W=0$:
+$R_{\text{dist}}(s)=-\|s-s^{*}\|.$
+However, $s_{W}$ may differ significantly from $s^{*}$, especially in degenerated situations where $W$ is large.
+Another reward signal could be to use the reference pattern $O^{*}$, which is known beforehand, like
+$R_{\text{pattern}}(s,W)=-\|O(s,W)-O^{*}\|.$
+Yet, even at the optimal position, $O(s_{W},W)$ may differ significantly from $O^{*}$ and we found that for complex patterns, the many local minima of $R_{\text{pattern}}$ make it hard for agents to learn good policies from that reward signal alone.
+Thus, we considered a combination of both rewards signals
+$R(s,W)=\begin{cases}R_{\text{pattern}}(s,W)+R_{\text{dist}}(s),\quad&\text{if }\|s-s^{*}\|\leq C\\ R_{\text{pattern}}(s,W),&\text{otherwise}\end{cases}.$
+This has the benefit that far away from $s^{*}$, and hence $s_{W}$, direction to the optimal region is enforced.
+Once being sufficiently close to $s^{*}$, the agents are purely guided by the pattern distance of the lens system at hand.
 IV Simulating Alignments
@@
 Here, we use Mitsuba3, a physically based rendering engine for forward and inverse light-transport simulation.
-This not only allows calculating light intensities $O_{W}(s)$ measured at the sensor (see Section IV-A) for a concrete position $s$, but also changing $s$ dynamically (see Section IV-B).
-All Mitsuba scenes consist of a sensor, $k$ many identical biconvex single lenses, and a light source.
+This not only allows calculating light intensities $O(s,W)$ measured at the sensor (see Section IV-A) for a concrete position $s$, but also changing $s$ dynamically (see Section IV-B).
+All Mitsuba scenes consist of a sensor, a lens system with four single lenses, and a light source.
+As one of our contribution is the open-source framework relign, we put high engineering effort in providing a realistic simulation environment and keeping the sim-to-real gap small.
 IV-A Sensor Outputs
-The Mitsuba scene emulates collimated light using a binary environment map with a centered rectangular light source.
+The Mitsuba scene emulates collimated light from a rectangular area emitter source and a specific reference pattern as binary bitmap.
+The light source is so far away that it can be considered infinity.
 Starting at the sensor, Mitsuba traces light rays backwards that pass through the lenses.
 We use an irradiance meter as sensor, which measures the incident power per unit area over a predefined shape.
-In our setup, the sensor shape is $0.5\times 0.5$ in Mitsuba space coordinates.
-To solve the high-dimensional problem of rendering $O_{W}(s)$ numerically, Mitsuba employs Monte Carlo integration, drawing $s_{c}$ many samples from a uniform distribution.
-As a consequence, the sensor output $O_{W}(s)$ is different when rendering the same position $s$ with different seeds.
-This leads to the probability density function of the POMDP as described in Section III-A.
+In our setup, the sensor shape is $2\times 2$ in Mitsuba space coordinates.
+To solve the high-dimensional problem of rendering $O(s,W)$ numerically, Mitsuba employs Monte Carlo integration, drawing $s_{c}$ many samples from a uniform distribution.
+As a consequence, the sensor output $O(s,W)$ is different when rendering the same position $s$ with different seeds.
+This leads to the probability density function of the POMDP as described in Section III.
 An example of the resulting sensor output is shown in Figure 1.
@@
 IV-B Position Changes
-Without loss of generality, we assume that the set of all possible states is the unit interval $[0,1]^{6}$.
-We place $k$ many lenses in the lens system $L$ in a way that $s^{*}:=(0.5,\ldots,0.5)\in\mathbb{R}^{6}$ is the optimal position.
+Without loss of generality, we assume that the set of all possible states is the unit interval $0,1]^{6}$.
+We place our lens system $L$ in a way that $s^{*}:=(0.5,\ldots,0.5)\in\mathbb{R}^{6}$ is the optimal position.
 Upon initialization, we sample $W_{\mathrm{L}}$ from a normal distribution and reposition the single lenses within $L$ accordingly.
@@
 The positioning of the lens system can be varied by setting an action $a\in\mathbb{R}^{6}$ leading to an update $s^{\prime}=s+W_{\mathrm{dist}}\cdot a$.
-To ensure that $s^{\prime}$ stays within $[0,1]^{6}$, each coordinate of $s+W_{\mathrm{dist}}\cdot a$ is clipped into $[0,1]$ before updating the Mitsuba scene.
-Subsequently, the sensor output $O_{W}(s^{\prime})$ is generated as described in Section IV-A.
+To ensure that $s^{\prime}$ stays within $0,1]^{6}$, each coordinate of $s+W_{\mathrm{dist}}\cdot a$ is clipped into $[0,1]$ before updating the Mitsuba scene.
+Subsequently, the sensor output $O(s^{\prime},W)$ is generated as described in Section IV-A.
 IV-C Generating the Reference Pattern
-To decide whether a state $s$ is in the optimality condition, the observation $O_{w}(s)$ has to be compared to a reference pattern $O^{*}$.
-This reference pattern only depends on the number of lenses $k$ in the lens system, not on the noise level of $W$.
-To generate $W$ in our synthetic setting, we use a perfectly aligned lens system where each of the $k$ many lenses are perfectly aligned as well, i.e., $W^{*}:=0$.
+To decide whether a state $s$ is in the optimality condition, the observation $O(s,W)$ has to be compared to a reference pattern $O^{*}$.
+This reference pattern only depends on the optical layout of the lens system, not on the noise level of $W$.
+To generate $W$ in our synthetic setting, we use a perfectly aligned lens system where each of the single lenses are perfectly aligned as well, i.e., $W^{*}:=0$.
 We then sample $1.000$ observations from $O_{W^{*}}(s^{*})$ and compute the pixel-wise mean image as an approximation for $O^{*}=\mathbb{E}[O_{W^{*}}(s^{*})]$.
-V Experiments
-V-A Benchmark Environments
+V Measuring the sim-to-real gap
+To emulate realistic manufacturing scenarios, we validate our simulation relign against an established optical design software and images from a real-world alignment station.
+For all the results presented in Section VI, we used the following hand-designed lens system was used.
+V-A Comparison to Optical Design Software
+To assess the applicability of the proposed workflow to a more realistic optical system, we designed a lens system consisting of four lenses in Code V [26] for monochromatic operation at 589 nm.
+The system is optimized for imaging objects at infinity onto the sensor plane and exhibits an effective focal length (EFL) of 2.895 mm, an f-number of 2.1, a field of view (FOV) of 84.1°, and a half exit-pupil angle of 13.45°.
+The lens system has a mechanical length of 20.59 mm and a maximum diameter of 9.8 mm.
+The design introduces a minimal distortion of –27.2% and throughout the entire field of view, the relative illumination does not fall below 60%.
+Figure 3 shows an optical diagram of the lens system including light rays of different wave lengths.
+Figure 3: Optical diagram of lens system including light rays in Code V.
+For integration into relign, each lens has to be exported individually from Code V and converted into polygonal meshes using any CAD tool (e.g., SolidWorks).
+Exporting the elements separately, rather than the lens system as a whole, ensures that the components remain independently addressable within the simulation environment.
+To enforce the defined field of view, we introduced an additional mechanical mount and an aperture stop directly behind the first lens.
+We modeled illumination by a rectangular area emitter providing diffuse emission with arbitrary bitmap textures.
+The emitter dimensions and distance were chosen such that the system operates within the optical field of view and the incident rays approximate collimated illumination, consistent with the design specifications.
+Simulation accuracy was validated using a chessboard test pattern.
+Subpixel-accurate corner positions were extracted from both the Code V and Mitsuba sensor outputs.
+The mean positional deviation amounted to approximately 1.2 pixels in both x- and y-directions, indicating close agreement between design and simulation as can be seen in Figure 4.
+(a) relign (ours)
+(b) Code V
+(c) Corner detection differences
+Figure 4: Comparison of point matching on a chessboard grid: relign (ours) vs. Code V.
+V-B Comparison to Real Active Alignment Situation
+Established or publicly documented alignment methods for complex multi-lens systems are rare hindering a realistic and unbiased comparison of different active alignment methods.
+This is mostly due to strong confidentiality constraints surrounding industrial active alignment, which typically cover algorithms, optical designs, hardware, and in some cases even the alignment targets.
+Consequently, we designed along given the optical parameters of a real-world camera system an optical design that approximates the performance.
+We then compare the output of relign to a real sensor output of $1920\times 1280$ pixel obtained through an alignment station.
+The visual comparison is shown in Figure 5 demonstrating that relign closely approximates a realistic scenario, with only a small deviation from the real-world camera.
+(a) relign (ours)
+(b) Real image from production
+(c) Difference relign (red) and production (green)
+Figure 5: Comparison of a realistic alignment target: relign (ours) vs. a real alignment station.
+VI Experiments
+We conduct a benchmark of different active alignment methods in this section to showcase how relign can be used. Throughout, we use the optical layout introduced in Section V.
 When rendering images, one typically faces a trade-off between computational efficiency and image quality.
 To reduce overall RL training time, we prioritize fast image generation at the cost of lower pixel resolution and increased sensor variance.
-Because of that, we only use $50\times 50$ pixels and $512$ samples per pixel for rendering.
-This also motivates the decision to use a simple rectangle as reference pattern, as more detailed patterns are not recognizable at this resolution.
-For the evaluation, we focus on six distinct benchmark setups.
-Each setup considers a set of $k\in\{2,3\}$ single lenses with radii of curvature of $-1.0$ and $3.8$ and thickness of $0.2$.
-Each single lens can have none, low or high individual variances $W_{L}$.
-The first three components of $W_{L}$ represent translation offsets along the $x$, $y$ and $z$ axes and are sampled from a normal distribution $(W_{L})_{i}\sim N_{0,6.25}$ for low and $(W_{L})_{i}\sim N_{0,12.5}$ for high variances with $i\in\{1,2,3\}$.
-The remaining two represent rotation offsets along the $x$ and $y$ axes, sampled as $(W_{L})_{j}\sim N_{0,0.00625}$ for low and $(W_{L})_{j}\sim N_{0,0.0125}$ for high variances with $j\in\{4,5\}$.
+Because of that, we only use $200\times 200$ pixels and $64$ samples per pixel for rendering.
+Accurate and comparable benchmarks for active alignments are hard to find, not only due to privacy constraints of companies maintaining the algorithms and the optical designs as part of their core know-how, but also due to the fact that optical products are very diverse and have different requirements.
+Already small differences in the required optical performance or the assumed variances of the optical system can lead to completely different alignment strategies and performances.
+Typical expert-designed alignment algorithms rely on hand-crafted features extracted from $O(s,W)$ that need to be optimized simultaneously, often via curve fitting [27] or expensive area scans.
+Already for use cases with two degrees of freedoms, dozens of steps may be required [27].
+A comparison of RL algorithms with classic methods on real systems thus requires a careful design to make them truly unbiased and insightful and hence this is beyond the scope of this paper.
+Nevertheless, we are convinced that our framework eases the effort to design fair and direct comparison in the future.
+Instead, we decided to benchmark against state-of-the art optimizers to solve Equation (1) actively and ensure that all algorithms are confronted with the exact same problem instances.
+VI-A Benchmark Environments
+For the evaluation, we focus on three distinct benchmark setups.
+Each setup considers the same lens system, but with either none, low or high individual variances $W_{L}$ for the last two lenses.
+The first three components of $W_{L}$ represent translation offsets along the $x$, $y$ and $z$ axes and are sampled from a normal distribution $(W_{L})_{i}\sim N_{0,1.25\cdot 10^{-4}}$ for low and $(W_{L})_{i}\sim N_{0,2.5\cdot 10^{-4}}$ for high variances with $i\in\{1,2,3\}$.
+The remaining two represent rotation offsets along the $x$ and $y$ axes, sampled as $(W_{L})_{j}\sim N_{0,0.375}$ for low and $(W_{L})_{j}\sim N_{0,0.75}$ for high variances with $j\in\{4,5\}$.
+We denote the underlying distributions as $\rho_{N}$ for $N=0$, $N=0.25$ and $N=0.5$.
 These variances account for potential manufacturing imperfections within a certain tolerance bound.
-To create a challenging alignment situation, they are intentionally set very high (see Figure 2 for an extreme case).
+To create a challenging alignment situation, they are intentionally set very high.
 Furthermore, rotation around $z$ is considered redundant, as in a perfectly aligned scenario, rotation around the $z$ axis has no effect.
 Note that $O^{*}$ does not change for setups including object variances.
-V-B Algorithms
-Expert-designed alignment algorithms typically rely on hand-crafted features extracted from $O_{W}(s)$ that need to be optimized simultaneously.
-These features as well as their optimality region highly depend on the exact use case.
-To have a fair and generic comparison of our RL approach with conventional learning methods, we decided to benchmark against state-of-the art black-box optimizers to solve Equation (1) actively.
-For our study, we focused on three different models to solve the task of optical alignments:
-V-B1 Proximal Policy Optimization
+VI-A1 RL Algorithms
 Proximal policy optimization (PPO) is a policy-based RL algorithm that trains a stochastic policy in an on-policy way [28].
-To process the images, a CNN is used as policy network.
-While PPO is a versatile learning algorithm, many learning tasks are too complex to solve from scratch via RL.
-Instead of training an agent on a hard task from the start, curriculum learning structures the training by introducing simpler subtasks first, enabling the agent to build foundational skills before tackling more challenging scenarios.
-This typically leads to faster convergence, better generalization, and improved sample efficiency, especially in environments with sparse rewards or long-horizon decision-making problems [27].
-The same holds true when training an RL agent for a too low optimality threshold $\theta$: The agent fails to reach the optimality region, even when allowing long trajectory lengths before truncation.
-Thus, we apply a curriculum strategy where $\theta$ is decreased gradually in an exponential fashion, that is, starting from an initial threshold $\theta_{0}$, the threshold of the $i$-th task is $\theta_{i}=C\cdot\theta_{i-1}$ with $C<1$ (in our experiments, we set $C=0.9$).
-We denote the task as solved when the agent manages to reach the optimality region defined by $\theta_{i}$ within $20$ steps in five subsequent online evaluations.
-The maximal length of an episode $T$ is set to $100$.
-All models have been trained for $1.5\cdot 10^{6}$ global steps with an initial learning rate of $1e^{-4}$ which is decreased by a factor of $0.8$ if the reward stagnates and which is reset to the initial value if the task of the curriculum is solved.
-Figure 4 shows the evolution of the reward along the curriculum for $k=3$ lenses.
-All trainings are executed on NVIDIA H100 GPUs requiring roughly $0.32$ seconds per global step resulting in a total train time of approximately five days.
-V-B2 Bayesian Optimization
+We used the PPO implementation in Stable Baselines3 [29] and compared it with other well-tested State-of-the-art RL algorithms.
+Our analysis revealed PPO’s superior performance, leading us to adopt it as the exclusive RL approach for this study.
+To process image observations, we employed a CNN as policy network, that consists of a standard CNN architecture with three $3\times 3$ convolutional layers, each followed by ReLU activation and max-pooling.
+This is followed by a single fully connected layer holding $256$ extracted features for each image.
+As is common in previous work, the learning rate and the entropy coefficient are crucial hyperpameters of PPO.
+We experimented with different learning rates in $\{1e^{-2},1e^{-3},1e^{-4},1e^{-5}\}$ and entropy coefficients in $\{1e^{-2},1e^{-3},1e^{-4}\}$.
+We set the discount factor to $\gamma=0.9$ and used the reward function described in Section III with $C=0.1$.
+The best-performing models were obtained after $1.9\cdot 10^{6}$ global steps and learning rates of $10^{-3}$ and $10^{-4}$, dependent on $W_{L}$.
+Figure 6 shows the evolution of the reward during online evaluations.
+Counterintuitively, one would expect to see a much slower convergence for higher noise during training.
+But due to the high variance in high-noise scenarios, the globally defined threshold $\theta$ must be set at a higher level to ensure that even lower-quality lens system still meet the acceptance criteria.
+Otherwise, such lens systems would be classified as not acceptable.
+However, increasing the tolerance in this way inevitably leads to a reduction in overall manufacturing quality.
+All trainings are executed on NVIDIA L40S GPUs requiring roughly $0.05$ seconds per global step resulting in a total train time of approximately $28$ hours.
+The full details can be found in the accompanying code of this paper.
+VI-A2 Bayesian Optimization
 Bayesian Optimization (BO) is a strategy for minimizing functions $f$ that are expensive to evaluate by building a probabilistic model of the objective function [30, 31].
@@
 This approach is particularly useful when function evaluations are costly, as it finds optimal solutions with relatively few evaluations.
-Setting $f_{W}(s):=\|O_{W}(s)-O^{*}\|$, an alignment problem can be interpreted as a black-box optimization problem.
+Setting $f_{W}(s):=-R(s,W)$, an alignment problem can be interpreted as an optimization problem for the family of black box functions $f_{W}$.
 We evaluated BO algorithms using a different probabilistic model than implemented in scikit-optimize [32]: Gaussian Processes (BO-GP) [33] and Random Forests (BO-RF) [34].
@@
 Thus, we also tested the method proposed in [35] (TransferGP) that allows pre-training a Gaussian process on samples from problem instances $f_{W_{1}},\ldots,f_{W_{m}}$.
-V-B3 Random
+VI-A3 Random
 As a baseline, we implemented an algorithm that samples uniformly at random from the alignment space.
-V-C Results
-For evaluation, each algorithm was executed on $100$ different environments for each of the six benchmark situations as described in Section V-A.
+VI-B Results
+For evaluation, each algorithm was executed on $100$ different environments for each of the benchmark situation described in Section VI-A.
 We compared approaches that operate without any a priori information and require no domain knowledge.
@@
 To address this imbalance, we reduced the search space for baseline algorithms to approximately eight percent of the search space used for RL.
-Figure 5 shows the RL-based method trained with $R_{\mathrm{pot}}$ surpases in all scenarios all other algorithms in terms of convergence speed.
-Moreover, independent of the number of lenses and the noise level, RL-PPO reaches the optimality region in under ten steps.
+The root mean squared error (RMSE), computed over all pixels, was used as the performance metric.
+Figure 7 shows that the RL-based method surpasses in all scenarios all other algorithms in terms of convergence speed.
+Moreover, independent of the noise level, RL-PPO reaches its optimum in under ten steps.
 As expected, the benchmarks with larger noise levels converge at a higher error, because due to the lens variances, even the best alignment cannot reach the minimal render variance error.
-In a direct comparison, the potential-based reward $R_{\mathrm{pot}}$ converges slighly faster than $R_{\mathrm{step}}$ to the optimal region (see Figure 6).
-Figure 4: Achieved reward $R_{\mathrm{pot}}$ during an online evaluation for $k=3$ lenses and different noise levels while optimality threshold $\theta$ is adjusted according to the curriculum.
+Figure 6: Evolution of rewards during training for different noise levels.
 Only considering the computation time, RL-PPO requires a constant 25 ms and BO-RF 95 ms per alignment step.
-In contrast, BO-GP becomes increasingly time consuming as the number of steps increases.
+In contrast, BO-GP becomes increasingly time-consuming as the number of steps increases.
 For 20 steps, each step takes an average of 100 ms, while for 50 steps the time per step increases to 178 ms.
@@
 Already when trained on $m=10$ instances with $100$ samples each, their processing time per step takes several minutes while their performance equals almost the performance of vanilla BO-GP.
-Due to their impractical computation time, we have not included pretrained GP models in our benchmark.
-Figure 5: Comparison of alignment algorithms.
-Figure 6: Evaluation of RMSE for the best PPO models.
-The upper plot compares two reward functions for a fixed number of single lenses (three).
-The lower plot compares two different lens systems using the best reward function.
-VI Conclusion
+Due to their impractical computation time, we have not included pre-trained GP models in our benchmark.
+Figure 7: Comparison of baseline alignment algorithms against RL-PPO for different noise levels.
+Figure 8: Evaluation of RMSE for the best PPO models and different noise levels.
+VII Conclusion
 This work introduced an RL approach for active alignments of optical components.
@@
 Future research could focus on improving sample efficiency, integrating domain adaptation techniques, and extending RL-based alignment to real-world hardware implementations.
-ACKNOWLEDGMENT
-We thank Marco F. Huber and our colleagues from the Institute of Machine Vision for their careful proofreading and constructive feedback on the manuscript.